OpenStudy (anonymous):

Trigonometric Equations: Solve each of the equations for 0≤x≤2π. Isolate the variable, sketch a reference triangle, find all the solutions within the given domain, 0≤x≤2π. Remember to double the domain when solving for a double angle. Use trig identities to rewrite the trig functions. sin x=-1/2

OpenStudy (anonymous):

@zepdrix HELP ME PLEASE

zepdrix (zepdrix):

They want you to draw a triangle for this? 0_o

zepdrix (zepdrix):

I guess we could draw a triangle or two like this. The sine function is negative in quadrants 3 and 4. \[\Large\rm \sin x=\frac{opposite}{hypotenuse}=\frac{-1}{2}\]

zepdrix (zepdrix):

Use the Pythagorean Theorem to find this missing side. I'm not sure how your teacher wants you to solve this. You can either relate it back to either the 45/45/90 or 30/60/90 triangle. Or you can use the inverse sine function.

OpenStudy (anonymous):

@zepdrix thanks for the help, i wasnt sure how to answer it either